# Positivity is undecidable in tensor products of free algebras

### 09 Dec 2023 v1

@@ 3.1. Dealing with infinite operator norm
[…]
Here sgn denotes the function $\mathbb{R}\to\{-1,1\}$ sending elements in $[0,+\infty)$ to $1$ and elements in $(-\infty,0)$ to $-1$.
@@
[…]
@@
where the last inequality follows from the conditions $(i)-(iv)$. By condition $(v)$, $\lVert(A-\tfrac{A+A^{*}}{2})\xi\rVert=\lVert\tfrac{A-A^{*}}{2}\xi\rVert\leq\tfrac{\epsilon}{2}$, so $\lVert(\widetilde{A}-A)\xi\rVert\leq\lVert(\widetilde{A}-\tfrac{A+A^{*}}{2})\xi\rVert+\lVert(A-\tfrac{A+A^{*}}{2})\xi\rVert\leq\tfrac{3}{2}\epsilon\leq 2\epsilon$. Similarly we have $\lVert(\widetilde{A}-A^{*})\xi\rVert\leq 2\epsilon$. ∎
@@
Let $\mathcal{X}$ be a finite set. There is a computable mapping $\mathcal{W}$ from *$\mathbb{Q}^{*}\langle\mathcal{X}\rangle$ to finite subsets of *$\mathbb{Q}^{*}\langle\mathcal{X}\rangle$ such that
@@
    $\mathcal{W}\big{(}\mathbb{Q}^{*}\langle\mathcal{X}\rangle\big{)}\subseteq\ker q$ where $q:\mathbb{C}^{*}\langle\mathcal{X}\rangle\to\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ is the quotient map, and
@@
    for any tracial state $\tau$ on *$\mathbb{C}^{*}\langle\mathcal{X}\rangle$, there is a tracial state $\widehat{\tau}$ on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ such that for all $f\in\mathbb{Q}^{*}\langle\mathcal{X}\rangle$ and $\epsilon\geq 0$, if $\tau$ is a tracial $(\epsilon,\mathcal{W}(f))$-state, then
@@
(so in particular, $\mathcal{W}(1)=\emptyset$). If $f\in\mathbb{Q}^{*}\langle\mathcal{X}\rangle$ has support $u_{1},\ldots,u_{\ell}$, or in other words $f=\sum_{i=1}^{\ell}\lambda_{i}u_{i}\in\mathbb{Q}^{*}\langle\mathcal{X}\rangle$ for distinct ***-monomials $u_{1},\ldots,u_{\ell}$ and non-zero coefficients $\lambda_{1},\ldots,\lambda_{\ell}$, we define $\mathcal{W}(f):=\bigcup_{i=1}^{\ell}\mathcal{W}(u_{i})$. Clearly $\mathcal{W}$ is computable. If $u$ is a monomial, then $q(s^{*}s)=0$ for all $s\in\mathcal{W}(u)$, so the same is true for $\mathcal{W}(f)$.
Let $\big{(}\pi:\mathbb{C}^{*}\langle\mathcal{X}\rangle\rightarrow\mathcal{B}(\mathcal{H}),v\big{)}$ be the GNS representation of $\tau$. Note that $T\mapsto\langle v,Tv\rangle$ extends to a tracial state on $\pi(\mathbb{C}^{*}\langle\mathcal{X}\rangle)^{\prime\prime}$. For every $x\in\mathcal{X}$, let $\widehat{\pi}(x):=\text{sgn}\big{(}\tfrac{\pi(x)+\pi(x)^{*}}{2}\big{)}$. By Lemma 3.10, $\widehat{\pi}(x)$ is a unitary involution in $C^{*}\langle\pi(x)\rangle^{\prime\prime}$ for every $x\in\mathcal{X}$, so $\widehat{\pi}$ induces a ***-representation of $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ on $\mathcal{H}$ and $\widehat{\tau}(a):=\langle v,\widehat{\pi}(a)v\rangle$ defines a tracial state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$.
@@
$\big{|}\lVert f\rVert_{\tau}-\lVert f\rVert_{\widehat{\tau}}\big{|}$ $=\big{|}\lVert\widehat{\pi}(f)v\rVert-\lVert\pi(f)v\rVert\big{|}\leq\lVert\big{(}\widehat{\pi}(f)-\pi(f)\big{)}v\rVert$
@@
Let $\mathcal{X}$ be a finite set, and let $\mathcal{W}$ be the map from Lemma 3.11. Suppose $\varphi$ is an $(\epsilon,\mathcal{X})$-synchronous state on *$\mathbb{C}^{*}\langle\mathcal{X}\rangle\otimes\mathbb{C}^{*}\langle\mathcal{X}\rangle$ for some $\epsilon\geq 0$, and $\varphi(s^{*}s\otimes 1)\leq\epsilon,\varphi(1\otimes s^{*}s)\leq\epsilon$ for all $s\in\bigcup_{x\in\mathcal{X}}\mathcal{W}(x)$. Then there is a $(25\epsilon,\mathcal{X})$-synchronous state $\widehat{\varphi}$ on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ such that for all $f\in\mathbb{Q}^{*}\langle\mathcal{X}\rangle$, if $\varphi(s^{*}s\otimes 1)\leq\epsilon$ for all $s\in\mathcal{W}(f)$, then
@@
Let $(\pi:\mathbb{C}^{*}\langle\mathcal{X}\rangle\otimes\mathbb{C}^{*}\langle\mathcal{X}\rangle\rightarrow\mathcal{B}(\mathcal{H}),\psi)$ be the GNS representation of $\varphi$. For every $x\in\mathcal{X}$, let $\widehat{\pi}(x\otimes 1):=\text{sgn}\big{(}\tfrac{\pi(x\otimes 1)+\pi(x\otimes 1)^{*}}{2}\big{)}$ and $\widehat{\pi}(1\otimes x):=\text{sgn}\big{(}\tfrac{\pi(1\otimes x)+\pi(1\otimes x)^{*}}{2}\big{)}$. Then by Lemma 3.10, for every $x\in\mathcal{X}$, $\widehat{\pi}(x\otimes 1)$ and $\widehat{\pi}(1\otimes x)$ are unitary involutions in $C^{*}\langle\pi(x\otimes 1)\rangle^{\prime\prime}$ and $C^{*}\langle\pi(1\otimes x)\rangle^{\prime\prime}$ respectively. So for every $x,x^{\prime}\in\mathcal{X}$, $[\pi(x\otimes 1),\pi(1\otimes x^{\prime})]=0$ implies $[\widehat{\pi}(x\otimes 1),\widehat{\pi}(1\otimes x^{\prime})]=0$, and hence $\widehat{\pi}$ induces a ***-representation of $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ on $\mathcal{H}$ and $\widehat{\varphi}(a):=\langle\psi,\widehat{\pi}(a)\psi\rangle$ defines a state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$. By the assumption, $\varphi(s^{*}s\otimes 1)\leq\epsilon$ and $\varphi(1\otimes s^{*}s)\leq\epsilon$ for all $s\in\bigcup_{x\in\mathcal{X}}\mathcal{W}(x)$. This implies $\lVert\pi(s\otimes 1)\psi\rVert\leq\sqrt{\epsilon}$ and $\lVert\pi(1\otimes s)\psi\rVert\leq\sqrt{\epsilon}$ for all $s\in\bigcup_{x\in\mathcal{X}}\mathcal{W}(x)$, and hence by Lemma 3.10, $\lVert\big{(}\widehat{\pi}(x\otimes 1)-\pi(x\otimes 1)\big{)}\psi\rVert\leq 2\sqrt{\epsilon}$ and $\lVert\big{(}\widehat{\pi}(1\otimes x)-\pi(1\otimes x)\big{)}\psi\rVert\leq 2\sqrt{\epsilon}$ for all $x\in\mathcal{X}$. It follows that
@@ 4. Construction of a suitable family of algebras
We begin the proofs of Theorems 1.1 and 1.3, with the case $\mathcal{A}=\mathbb{Q}\mathbb{Z}_{2}^{*N}$. To explain the idea of the proof, suppose that $\mathcal{S}\subseteq\mathbb{N}$ is an RE set, and that we can find an integer $N$, a self-adjoint element $w\in\mathbb{Q}\mathbb{Z}_{2}^{*N}$, and a computable family of finitely presented algebras $\mathcal{A}_{\mathcal{S}}(m)=\mathbb{C}^{*}\langle\mathbb{C}\mathbb{Z}_{2}^{*N}:\mathcal{R}_{m}\rangle$ such that $\mathcal{A}_{\mathcal{S}}(m)$ has a tracial state $\tau$ with $\tau(w^{2})>0$ if and only if $m\in\mathcal{S}$. If $\alpha(m):=\sum_{r\in\mathcal{R}_{m}}r^{*}r-w^{2}$, then a tracial state $\tau$ on $\mathcal{A}_{\mathcal{S}}(m)$ with $\tau(w^{2})>0$ pulls back to a tracial state on $\mathbb{C}\mathbb{Z}_{2}^{*N}$ with $\tau(\alpha(m))<0$. Unfortunately this idea runs into a stability problem: it is not necessarily the case that every tracial state on $\mathbb{C}\mathbb{Z}_{2}^{*N}$ with $\tau(\alpha(m))<0$ comes from a tracial state on $\mathcal{A}_{\mathcal{S}}(m)$. Indeed, it might be possible to find a tracial state $\tau$ for which $\tau\left(\sum_{r\in\mathcal{R}_{m}}r^{*}r\right)$ is small, but $\tau(w^{2})$ is large.
The main technical idea of this paper is that we can overcome the stability problem with a carefully chosen algebra $\mathcal{A}_{\mathcal{S}}(m)$. Specifically, we start with the algebra generated by a family of positive contractions $\{\widetilde{P}_{0},\widetilde{P}_{1},\ldots\}$ and a family of unitary involutions $\{\widetilde{X}_{0},\widetilde{X}_{1},\ldots\}$ satisfying the relations
@@
for all $n$ such that $M_{\mathcal{S}}$ does not halt on input $m$ after $n$ steps, where $M_{\mathcal{S}}$ is a Turing machine recognizing $\mathcal{S}$. This algebra is not finitely presented, so the purpose of this section is to show that this algebra can be embedded in a finitely-presented algebra $\mathcal{A}_{\mathcal{S}}(m)$, in such a way that the size of relation (4.1) is polynomial in $n$. Relation (4.1) implies that $\tau(\widetilde{P}_{n})=\tfrac{1}{2}\tau(\widetilde{P}_{n+1})$ for any tracial state $\tau$ on $\mathcal{A}_{\mathcal{S}}(m)$, and in the next section we use this fact to overcome the stability problem and finish the proofs of Theorems 1.1 and 1.3 (for the case $\mathcal{A}=\mathbb{Q}\mathbb{Z}_{2}^{*N}$) using the strategy outlined in the previous paragraph.
@@
Let $\mathcal{S}\subset\mathbb{N}$ be an RE set, and let $M_{\mathcal{S}}$ be a Turing machine that recognizes $\mathcal{S}$. Then there is a finite set $\mathcal{X}$, two sequences $\{\widetilde{P}_{0},\widetilde{P}_{1},...\}$ and $\{\widetilde{X}_{0},\widetilde{X}_{1},...\}$ in *$\mathbb{Q}^{*}\langle\mathcal{X}\rangle$, and a computable mapping $m\mapsto\mathcal{R}_{m}$ from $\mathbb{N}$ to finite subsets of $\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}$, such that
@@
    for all $n\geq 0$, $\widetilde{P}_{n}$ is a hermitian square in $\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}$ with $\lVert\widetilde{P}_{n}\rVert_{1}=1$ and $\lVert\widetilde{P}_{n}\rVert_{1,1}=3(n+1)/2$, and $\widetilde{X}_{n}$ is a monomial in *$\mathbb{Q}^{*}\langle\mathcal{X}\rangle$ of degree $14n+$ 1 such that $\widetilde{X}_{n}^{2}=1$ in $\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}$,
@@
    there are constants $C>0,k\in\mathbb{N}$ such that for every $m\in\mathbb{N}$ and $n\geq 0$, if $M_{\mathcal{S}}$ hasn’t halted after $n$ steps on input $m$, then $\widetilde{P}_{n}+\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}-\widetilde{P}_{n+1}$ is trivial in $\mathcal{A}_{\mathcal{S}}(m):=\mathbb{C}^{*}\langle\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}:\mathcal{R}_{m}\rangle$ and has an $\mathcal{R}_{m}$-decomposition in $\mathbb{C}\mathbb{Z}_{2}^{\mathcal{X}}$ of size $\leq C\big{(}(n+1)m\big{)}^{k}$,
@@
    if $m\in\mathcal{S}$, then there is a tracial state $\tau$ on $\mathcal{A}_{\mathcal{S}}(m)$ such that $\tau(\widetilde{P}_{0}^{2})>0$, and
@@
In part (a), $\lVert\cdot\rVert_{1}$ and $\lVert\cdot\rVert_{1,1}$ denote the 1-norm and Sobolev seminorm on *$\mathbb{C}^{*}\langle\mathcal{X}\rangle$ defined in the previous section. Since monomials in *$\mathbb{C}^{*}\langle\mathcal{X}\rangle$ are unitary in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$, $\lVert a\rVert_{\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}}\leq\lVert a\rVert_{1}$ for all $a\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$. So part (a) implies that $\widetilde{P}_{n}$ is a contraction in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$, i.e. $\lVert\widetilde{P}_{n}\rVert_{\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}}\leq 1$, and $\widetilde{X}_{n}$ is a unitary involution in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ for all $n\geq 0$.
For the rest of this section, we fix an RE set $\mathcal{S}\subset\mathbb{N}$, and a deterministic Turing machine $M_{\mathcal{S}}$ that recognizes $\mathcal{S}$. Define $h:\mathbb{Z}\rightarrow\mathbb{N}\cup\{\infty\}$ by
$h(m)=\begin{cases}n&m\in\mathcal{S}\text{ and }M_{\mathcal{S}}\text{ halts at the }n\text{th step on the input }m\\ \infty&m\notin\mathcal{S}\end{cases},$
so $h|_{\mathbb{N}}$ is the running time of $M_{\mathcal{S}}$. Recall that in our convention, all Turing machines take at least one step, so $h(m)\geq 1$ for all $m\in\mathbb{Z}$. To prove Theorem 4.1, we start in Section 4.1 by encoding the Turing machine $M_{\mathcal{S}}$ in certain group relations. We finish the proof by constructing $\mathcal{A}_{\mathcal{S}}(m)$ in Section 4.2.

### 18 Mar 2026 v2

@@ 3.1. Dealing with infinite operator norm
For any $A\in\mathcal{B}(\mathcal{H})$, the operator $\widetilde{A}:=\text{sgn}\big(\frac{A+A^{*}}{2}\big)$ is a unitary involution in $C^{*}\langle A\rangle^{\prime\prime}$. If in addition, $\xi$ is a vector in $\mathcal{H}$ satisfying (i) $\lVert(A^{2}-\mathbbm{1}_{\mathcal{H}})\xi\rVert\leq\epsilon$, (ii) $\lVert((A^{*})^{2}-\mathbbm{1}_{\mathcal{H}})\xi\rVert\leq\epsilon$, (iii) $\lVert(AA^{*}-\mathbbm{1}_{\mathcal{H}})\xi\rVert\leq\epsilon$, (iv) $\lVert(A^{*}A-\mathbbm{1}_{\mathcal{H}})\xi\rVert\leq\epsilon$, and (v) $\lVert(A-A^{*})\xi\rVert\leq\epsilon$ for some $\epsilon\geq 0$, then $\lVert(A-\widetilde{A})\xi\rVert\leq 2\epsilon$ and $\lVert(A^{*}-\widetilde{A})\xi\rVert\leq 2\epsilon$.
Here sgn denotes the function $\mathbb{R}\to\{-1,1\}$ sending elements in $0,+\infty)$ to $1$ and elements in $(-\infty,0)$ to 1-1.
@@
The spectrum of $\tfrac{A+A^{*}}{2}$ is in $\mathbb{R}$, so by Borel functional calculus, $\widetilde{A}:=\text{sgn}(\tfrac{A+A^{*}}{2})$ is a unitary involution and is in $C^{*}\langle A\rangle^{\prime\prime}$. Since sgn and $\lvert\cdotp\rvert$ are measurable functions on $\mathbb{R}$ satisfying $(1+\lvert a\rvert)(\text{sgn}(a)-a)=\text{sgn}(a)(1-a^{2})$ for all $a\in\mathbb{R}$, we obtain that
@@
where the last inequality follows from the conditions (i)(iv)(i)-(iv). By condition $(v)$, $\lVert(A-\tfrac{A+A^{*}}{2})\xi\rVert=\lVert\tfrac{A-A^{*}}{2}\xi\rVert\leq\tfrac{\epsilon}{2}$, so $\lVert(\widetilde{A}-A)\xi\rVert\leq\lVert(\widetilde{A}-\tfrac{A+A^{*}}{2})\xi\rVert+\lVert(A-\tfrac{A+A^{*}}{2})\xi\rVert\leq\tfrac{3}{2}\epsilon\leq 2\epsilon$. Similarly we have $\lVert(\widetilde{A}-A^{*})\xi\rVert\leq 2\epsilon$. ∎
@@
Let $\mathcal{X}$ be a finite set. There is a computable mapping $\mathcal{W}$ from $\mathbb{Q}^{*}\langle\mathcal{X}\rangle$ to finite subsets of $\mathbb{Q}^{*}\langle\mathcal{X}\rangle$ such that
@@
    $\mathcal{W}(f)\subseteq\ker q$ for all $f\in\mathbb{Q}^{*}\langle\mathcal{X}\rangle$, where $q:\mathbb{C}^{*}\langle\mathcal{X}\rangle\to\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ is the quotient map, and
@@
    for any tracial state $\tau$ on $\mathbb{C}^{*}\langle\mathcal{X}\rangle$, there is a tracial state $\widehat{\tau}$ on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ such that for all $f\in\mathbb{Q}^{*}\langle\mathcal{X}\rangle$ and $\epsilon\geq 0$, if $\tau$ is a tracial $(\epsilon,\mathcal{W}(f))$-state, then
@@
(so in particular, $\mathcal{W}(1)=\emptyset$). If $f\in\mathbb{Q}^{*}\langle\mathcal{X}\rangle$ has support $u_{1},\ldots,u_{\ell}$, or in other words $f=\sum_{i=1}^{\ell}\lambda_{i}u_{i}\in\mathbb{Q}^{*}\langle\mathcal{X}\rangle$ for distinct *-monomials $u_{1},\ldots,u_{\ell}$ and non-zero coefficients $\lambda_{1},\ldots,\lambda_{\ell}$, we define $\mathcal{W}(f):=\bigcup_{i=1}^{\ell}\mathcal{W}(u_{i})$. Clearly $\mathcal{W}$ is computable. If $u$ is a monomial, then $q(s^{*}s)=0$ for all $s\in\mathcal{W}(u)$, so the same is true for $\mathcal{W}(f)$.
Let $\big(\pi:\mathbb{C}^{*}\langle\mathcal{X}\rangle\rightarrow\mathcal{B}(\mathcal{H}),v\big)$ be the GNS representation of $\tau$. Note that $T\mapsto\langle v,Tv\rangle$ extends to a tracial state on $\pi(\mathbb{C}^{*}\langle\mathcal{X}\rangle)^{\prime\prime}$. For every $x\in\mathcal{X}$, let $\widehat{\pi}(x):=\text{sgn}\big(\tfrac{\pi(x)+\pi(x)^{*}}{2}\big)$. By Lemma 3.10, $\widehat{\pi}(x)$ is a unitary involution in $C^{*}\langle\pi(x)\rangle^{\prime\prime}$ for every $x\in\mathcal{X}$, so $\widehat{\pi}$ induces a *-representation of $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ on $\mathcal{H}$ and $\widehat{\tau}(a):=\langle v,\widehat{\pi}(a)v\rangle$ defines a tracial state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$.
@@
[…]
@@
Let $\mathcal{X}$ be a finite set, and let $\mathcal{W}$ be the map from Lemma 3.11. Suppose $\varphi$ is an $(\epsilon,\mathcal{X})$-synchronous state on $\mathbb{C}^{*}\langle\mathcal{X}\rangle\otimes\mathbb{C}^{*}\langle\mathcal{X}\rangle$ for some $\epsilon\geq 0$, and $\varphi(s^{*}s\otimes 1)\leq\epsilon,\varphi(1\otimes s^{*}s)\leq\epsilon$ for all $s\in\bigcup_{x\in\mathcal{X}}\mathcal{W}(x)$. Then there is a $(25\epsilon,\mathcal{X})$-synchronous state $\widehat{\varphi}$ on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ such that for all $f\in\mathbb{Q}^{*}\langle\mathcal{X}\rangle$, if $\varphi(s^{*}s\otimes 1)\leq\epsilon$ for all $s\in\mathcal{W}(f)$, then
@@
Let $(\pi:\mathbb{C}^{*}\langle\mathcal{X}\rangle\otimes\mathbb{C}^{*}\langle\mathcal{X}\rangle\rightarrow\mathcal{B}(\mathcal{H}),\psi)$ be the GNS representation of $\varphi$. For every $x\in\mathcal{X}$, let $\widehat{\pi}(x\otimes 1):=\text{sgn}\big(\tfrac{\pi(x\otimes 1)+\pi(x\otimes 1)^{*}}{2}\big)$ and $\widehat{\pi}(1\otimes x):=\text{sgn}\big(\tfrac{\pi(1\otimes x)+\pi(1\otimes x)^{*}}{2}\big)$. Then by Lemma 3.10, for every $x\in\mathcal{X}$, $\widehat{\pi}(x\otimes 1)$ and $\widehat{\pi}(1\otimes x)$ are unitary involutions in $C^{*}\langle\pi(x\otimes 1)\rangle^{\prime\prime}$ and $C^{*}\langle\pi(1\otimes x)\rangle^{\prime\prime}$ respectively. So for every $x,x^{\prime}\in\mathcal{X}$, $\pi(x\otimes 1),\pi(1\otimes x^{\prime})]=0$ implies $\widehat{\pi}(x\otimes 1),\widehat{\pi}(1\otimes x^{\prime})]=0$, and hence $\widehat{\pi}$ induces a *-representation of $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ on $\mathcal{H}$ and $\widehat{\varphi}(a):=\langle\psi,\widehat{\pi}(a)\psi\rangle$ defines a state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$. By the assumption, $\varphi(s^{*}s\otimes 1)\leq\epsilon$ and $\varphi(1\otimes s^{*}s)\leq\epsilon$ for all $s\in\bigcup_{x\in\mathcal{X}}\mathcal{W}(x)$. This implies $\lVert\pi(s\otimes 1)\psi\rVert\leq\sqrt{\epsilon}$ and $\lVert\pi(1\otimes s)\psi\rVert\leq\sqrt{\epsilon}$ for all $s\in\bigcup_{x\in\mathcal{X}}\mathcal{W}(x)$, and hence by Lemma 3.10, $\lVert\big(\widehat{\pi}(x\otimes 1)-\pi(x\otimes 1)\big)\psi\rVert\leq 2\sqrt{\epsilon}$ and $\lVert\big(\widehat{\pi}(1\otimes x)-\pi(1\otimes x)\big)\psi\rVert\leq 2\sqrt{\epsilon}$ for all $x\in\mathcal{X}$. It follows that
@@ 4. Construction of a suitable family of algebras
We begin the proofs of Theorems 1.1 and 1.3, with the case $\mathcal{A}=\mathbb{Q}\mathbb{Z}_{2}^{*N}$. To explain the idea of the proof, suppose that $\mathcal{L}\subseteq\mathbb{N}$ is an RE set, and that we can find an integer $N$, a self-adjoint element $w\in\mathbb{Q}\mathbb{Z}_{2}^{*N}$, and a computable family of finitely presented algebras $\mathcal{A}_{\mathcal{L}}(m)=\mathbb{C}^{*}\langle\mathbb{C}\mathbb{Z}_{2}^{*N}:\mathcal{R}_{m}\rangle$ such that $\mathcal{A}_{\mathcal{L}}(m)$ has a tracial state $\tau$ with $\tau(w^{2})>0$ if and only if $m\in\mathcal{L}$. If $\alpha(m):=\sum_{r\in\mathcal{R}_{m}}r^{*}r-w^{2}$, then a tracial state $\tau$ on $\mathcal{A}_{\mathcal{L}}(m)$ with $\tau(w^{2})>0$ pulls back to a tracial state on $\mathbb{C}\mathbb{Z}_{2}^{*N}$ with $\tau(\alpha(m))<0$. Unfortunately this idea runs into a stability problem: it is not necessarily the case that every tracial state on $\mathbb{C}\mathbb{Z}_{2}^{*N}$ with $\tau(\alpha(m))<0$ comes from a tracial state on $\mathcal{A}_{\mathcal{L}}(m)$. Indeed, it might be possible to find a tracial state $\tau$ for which $\tau\left(\sum_{r\in\mathcal{R}_{m}}r^{*}r\right)$ is small, but $\tau(w^{2})$ is large.
The main technical idea of this paper is that we can overcome the stability problem with a carefully chosen algebra $\mathcal{A}_{\mathcal{L}}(m)$. Specifically, we start with the algebra generated by a family of positive contractions $\{\widetilde{P}_{0},\widetilde{P}_{1},\ldots\}$ and a family of unitary involutions $\{\widetilde{X}_{0},\widetilde{X}_{1},\ldots\}$ satisfying the relations
@@
for all $n$ such that $M_{\mathcal{L}}$ does not halt on input $m$ after $n$ steps, where $M_{\mathcal{L}}$ is a Turing machine recognizing $\mathcal{L}$. This algebra is not finitely presented, so the purpose of this section is to show that this algebra can be embedded in a finitely presented algebra $\mathcal{A}_{\mathcal{L}}(m)$, in such a way that the size of relation (4.1) is polynomial in $n$. Relation (4.1) implies that $\tau(\widetilde{P}_{n})=\tfrac{1}{2}\tau(\widetilde{P}_{n+1})$ for any tracial state $\tau$ on $\mathcal{A}_{\mathcal{L}}(m)$, and in the next section we use this fact to overcome the stability problem and finish the proofs of Theorems 1.1 and 1.3 (for the case $\mathcal{A}=\mathbb{Q}\mathbb{Z}_{2}^{*N}$) using the strategy outlined in the previous paragraph.
@@
Let $\mathcal{L}\subset\mathbb{N}$ be an RE set, and let $M_{\mathcal{L}}$ be a Turing machine that recognizes $\mathcal{L}$. Then there is a finite set $\mathcal{X}$, two sequences $\{\widetilde{P}_{0},\widetilde{P}_{1},$...\} and $\{\widetilde{X}_{0},\widetilde{X}_{1},$...\} in $\mathbb{Q}^{*}\langle\mathcal{X}\rangle$, and a computable mapping $m\mapsto\mathcal{R}_{m}$ from $\mathbb{N}$ to finite subsets of $\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}$, such that
@@
    for all $n\geq 0$, $\widetilde{P}_{n}$ is a hermitian square in $\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}$ with $\lVert\widetilde{P}_{n}\rVert_{1}=1$ and $\lVert\widetilde{P}_{n}\rVert_{1,1}=3(n+1)/2$, and $\widetilde{X}_{n}$ is a monomial in $\mathbb{Q}^{*}\langle\mathcal{X}\rangle$ of degree $4n+1$ such that $\widetilde{X}_{n}^{2}=1$ in $\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}$,
@@
    there are constants $C>0,k\in\mathbb{N}$ such that for every $m\in\mathbb{N}$ and $n\geq 0$, if $M_{\mathcal{L}}$ hasn’t halted after $n$ steps on input $m$, then $\widetilde{P}_{n}+\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}-\widetilde{P}_{n+1}$ is trivial in $\mathcal{A}_{\mathcal{L}}(m):=\mathbb{C}^{*}\langle\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}:\mathcal{R}_{m}\rangle$ and has an $\mathcal{R}_{m}$-decomposition in $\mathbb{C}\mathbb{Z}_{2}^{\mathcal{X}}$ of size $\leq C\big((n+1)m\big)^{k}$,
@@
    if $m\in\mathcal{L}$, then there is a tracial state $\tau$ on $\mathcal{A}_{\mathcal{L}}(m)$ such that $\tau(\widetilde{P}_{0}^{2})>0$, and
@@
In part (a), $\lVert\cdot\rVert_{1}$ and $\lVert\cdot\rVert_{1,1}$ denote the 1-norm and Sobolev seminorm on $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ defined in the previous section. Since monomials in $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ are unitary in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$, $\lVert a\rVert_{\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}}\leq\lVert a\rVert_{1}$ for all $a\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$. So part (a) implies that $\widetilde{P}_{n}$ is a contraction in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$, i.e. $\lVert\widetilde{P}_{n}\rVert_{\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}}\leq 1$, and $\widetilde{X}_{n}$ is a unitary involution in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ for all $n\geq 0$.
For the rest of this section, we fix an RE set $\mathcal{L}\subset\mathbb{N}$, and a deterministic Turing machine $M_{\mathcal{L}}$ that recognizes $\mathcal{L}$. Define $h:\mathbb{Z}\rightarrow\mathbb{N}\cup\{\infty\}$ by
[…]
so $h|_{\mathbb{N}}$ is the running time of $M_{\mathcal{L}}$. Recall that in our convention, all Turing machines take at least one step, so $h(m)\geq 1$ for all $m\in\mathbb{Z}$. To prove Theorem 4.1, we start in Section 4.1 by encoding the Turing machine $M_{\mathcal{L}}$ in certain group relations. We finish the proof by constructing $\mathcal{A}_{\mathcal{L}}(m)$ in Section 4.2.